Convert the following linear equations from standard form to y-intercept form step by step.
1.a) 5x-y+3=0
b) x+7y-14=0
c)6x-3y+9=0
d)7x+2y-3=0
e) 2x+5y+5=0
f) 12x-30y+13=0

Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

Convert the following linear equations from standard form to y-intercept form step by step.
1.a) 5x-y+3=0
------
That means solve for y.
5x-y+3=0
5x+3 = y
y = 5x+3
=============
Do the others the same way.

Answer by Boreal(15235)

Isolate the y and then make it 1y by dividing everything in the equation, the x term and the constant by whatever the number is in front of the y (its coefficient).
a)here, it is easier to move the y and leave the others alone. y=5x+3. Add y to both sides
b)Here, move everything else but the y so 7y=-x+14; subtract the x and and 14 to both sides
divide everything by 7 to get y=(-1/7)x+2
c)move the y to the right to get 6x+9=3y. Then divide by 3 to get 2x+3=y
d)move the x and constant to get 2y=-7x+3 or y=-(7/2)x+(3/2)
e) same, 5y=-2x-5 and divide by 5 to get y=-(2/5)x-1
f)move the y then divide everything by 30--30y=12x+13 or y=(12/30)x+(13/30) and can reduce 12/30 to 2/5.